1) A saturated solution of copper (II) iodate in pure water has a copper ion concentration of 2.7X10^-3 M.

a) whar is the molar solubility of copper iodate in a 0.35 M Potassium iodate solution

B) what is the molar solubility of copper Iodate in 0.35 M copper nitrate solution

2)

a)consider the slightly soluble salt, silver chloride, Agcl. If you had 10.0 g of silver chloride, what volume of water would be required to completely dissolve this amount

b) what volume of 15 M NH3 would be required to dissolve the same quantity of AgCl completey
Answer #1

1 Ans.

The dissociation of Cu(IO3)2 is

Cu(IO_{3})_{2}\rightleftharpoons Cu^{2+}+ 2IO_{3}^{-}

The Ksp expression is

K_{sp}= [Cu^{2+}][IO_{3}^{-}]^{2} = x(2x)^{2}

where x is the molar solubility of Cu(IO3)2 and also the concentration of copper ion. The solution is a saturated solution. So, the concentration of Cu2+ is the same as the molar solubility.

K_{sp} = x(2x)^{2}

K_{sp} = 4x^{3}

Given that,

x = 2.7 \times 10^{-3}M

Therefore,

K_{sp} = 4\times (2.7\times 10^{-3})^{3}

K_{sp} = 7.87 \times 10^{-8}

Therefore, Ksp of Cu(IO3)2 is 7.87 x 10-8

a Ans.

Molarity of KIO3 = 0.35M

[IO3-] from KIO3 = 0.35M

Ksp of Cu(IO3)2 as written before is

K_{sp}= [Cu^{2+}][IO_{3}^{-}]^{2} = x(2x)^{2}

But, 0.35 has to be added to the concentration of IO3-

Therefore,

K_{sp} = x(2x+ 0.35)^{2}

The value of 2x is very negligibly small compared to 0.35M. So, 2x is neglected to simplify the calculation. Neglecting the 2x term will cause little or no difference.

K_{sp} = x(0.35)^{2}

Substitute the value of Ksp

7.87\times 10^{-8} = x(0.35)^{2}

7.87\times 10^{-8} = 0.1225x

\frac{7.87\times 10^{-8} }{0.1225}= x

x = 6.4 \times 10^{-7} mol/L

The molar solubility of copper iodate in 0.35M potassium iodate solution is 6.4 x 10-7mol/L.

b Ans.

Concentration of Cu2+ from copper nitrtate solution = 0.35M

Ksp of Cu(IO3)2 as written before is

K_{sp}= [Cu^{2+}][IO_{3}^{-}]^{2} = x(2x)^{2}

To the concentration of copper ion, 0.35M has to be added which is already present in the copper nitrate solution.

Therefore,

K_{sp} = (x+0.35)(2x)^{2}

K_{sp} = (0.35)(2x)^{2}

x, being very small compared to 0.35 is neglected to simplify calculation.

K_{sp} = 0.35\times 4x^{2}

K_{sp} = 1.40x^{2}

7.87\times 10^{-8} = 1.40x^{2}

\frac{7.87\times 10^{-8}}{1.40}= x^{2}

\sqrt{\frac{7.87\times 10^{-8}}{1.40}}= x

x = 2.4 \times 10^{-4}mol/L

Therefore, the molar solubility of copper iodate in 0.35M copper nitrate solution is 2.4 x 10-4 mol/L.

2 a Ans

The Ksp value of AgCl at 25 degree celcius = 1.8 x 10-10

AgCl\rightleftharpoons Ag^{+}+Cl^{-}

K_{sp}= [Ag^{+}][Cl^{-}]=x^{2}

1.8 \times 10^{-10} =x^{2}

\sqrt{1.8 \times 10^{-10} }=x

x= 1.34 \times 10^{-5}mol/L

Molar mass of AgCl = 143.32g/mol

143.32g of AgCl is 1 mol.

Therefore, 10.0g of AgCl is

= \frac{1}{143.32g/mol}\times 10.0g

= 0.06977 mol

Molar solubility, x= 1.34 \times 10^{-5}mol/L

1.34 x 10-5 mol of AgCl is soluble in 1 L

So, 0.0698 mol of AgCl will be soluble in

= \frac{1}{1.34\times 10^{-5} mol/L}\times 0.06977mol

= 5206.72L

Therefore, 5207L of water is required to completely dissolve 10.0g AgCl.

b Ans. Molarity of NH3 = 15M

AgCl(s)\rightleftharpoons Ag^{+}(aq)+ Cl^{-}(aq)

Ksp = 1.8 x 10-10

2NH_{3}(aq)+ Ag^{+}(aq)\rightleftharpoons [Ag(NH_{3})_{2}]^{+}(aq)

Kf = 1.6 x 107

The overall reaction is

AgCl(s)+2NH_{3}(aq) \rightleftharpoons [Ag(NH_{3})_{2}]^{+}(aq)+ Cl^{-}(aq)

The equilibrium constant of the overall reaction is

K = K_{sp}\times K_{f}

K = 1.8 \times 10^{-10} \times 1.6 \times 10^{7}

K = 2.88 \times 10^{-3}

The ICE table for the overall reaction is

AgCl(s)+2NH_{3}(aq) \rightleftharpoons [Ag(NH_{3})_{2}]^{+}(aq)+ Cl^{-}(aq)

AgCl(s) NH3(aq) [Ag(NH3)2]+ Cl-
I - 15M 0 0
C - -2x +x +x
E - 15-2x x x

K = \frac{[Ag(NH_{3})_{2}]^{+}][Cl^{-}]}{[NH_{3}]^{2}}

2.88 \times 10^{-3}= \frac{x^{2}}{(15-2x)^{2}}

Take square root on both sides,

\sqrt{2.88 \times 10^{-3}}= \frac{x}{15-2x}

0.05367= \frac{x}{15-2x}

0.05367(15-2x)= x

0.8051-0.1073x= x

0.8051= 1.1073x

x =\frac{0.8051}{1.1073}

x =0.7271 mol/L

0.7271 mol of AgCl is soluble in 1L ammonia

10.0g AgCl has 0.0698 mol (calculated in part a)

0.0698 mol will be soluble in

=\frac{1}{0.7271 mol/L}\times 0.0698mol

=0.096L

=96mL

Therefore, 0.096L of 15M ammonia is required.

(N.B: Substitute the exact values of the constants given to you to get the best result and follow the same approach to solve.)